OpenStudy (anonymous):

Choose the correct simplification of the expression (3x)^4. My answer is 12x^4 Am I right? :)

OpenStudy (nurali):

3^4=3*3*3*3=81

OpenStudy (anonymous):

^

OpenStudy (nurali):

so 81x^4

OpenStudy (anonymous):

Ok, Thanks

OpenStudy (nurali):

Anytime.